# AI-powered simulation-based inference of a genuinely spatial-stochastic model of early mouse embryogenesis Michael A. Ramirez-Sierra1,2, Thomas R. Sokolowski1*

### 23 Feb 2024 v1

@@ Discussion
The first cell-fate decision happens between E2.5 and E3.0 (from 8- to 32-cell stage): cells acquire TE or ICM identities. The second cell-fate decision happens at the ICM between E3.0 and E4.0 (from 32- to 128-cell stage). From E4.0 to 4.5, EPI and PRE populations spatially separate. While it is customary to define the blastocyst-formation period between E3.0 and E4.5 [93, 4], these boundaries are ultimately arbitrary as development occurs in a continuum and diverse experimental arrangements/conditions are in use between distinct labs. Moreover, ICM cells adopt their next identities asynchronously as the blastocyst forms [114, 14, 122]. Together with these aspects, it is also commonly accepted that cells are already coexpressing Nanog- and Gata6-related factors at around E2.75 [103, 104], plus Fgf4 expression is already perceptible at around E3.25 [93, 4]. For these reasons, our standard model simulations target a time-window of 48 hours (E2.75-E4.75); this range allows us to circumvent potential discrepancies among timing annotations and keep a temporally faithful description of the biological system under study.
@@
The core GRN motif is exclusively comprised by Nanog- and Gata6-related elements. To be more precise, all their directly related species and interactions. The rest of the full GRN is built around the core motif, thus consolidating the remaining elements and their collective effects on the dynamics of the two main players. Importantly, we have arranged all cell-scale reaction events into several groups as follows: summary of gene expression dynamics; promoter binding and unbinding; mRNA synthesis and degradation; protein synthesis and degradation; FGFR activation and inactivation; ERK activation and inactivation; NANOG phosphorylation and dephosphorylation. In that regard, we report all the particular interactions implemented by our simulator and their respective literature sources.
@@
We model TF cooperativity by expressly tuning the promoter unbinding rates. This rate tuning influences the promoter regulation model to mimic a Hill-function-like (nonlinear) transcriptional response. The usage of the Hill function is a staple of phenomenological modelling, however it is incompatible with mechanistic modelling; directly using a Hill equation as a reaction propensity function ignores the non-instantaneous (stochastic) nature of delays between biochemical events and introduces several other simulation artifacts [49, 22, 17]. We use the concepts of a half-saturation constant and a cooperativity coefficient to perform promoter unbinding rate tuning; accordingly, both quantities are incorporated into elementary reactions to describe promoter unbinding dynamics. This half-saturation constant (hact or hrep depending on TF role) is a free model parameter which dictates the threshold of TF copies needed for reaching 50% of negative or positive gene transcriptional control; consequently, each gene-TF pair requires its own separate half-saturation threshold. This cooperativity coefficient (kcoop) is an auxiliary variable which adjusts the strength of mutual influence among TF copies; we arbitrarily defined it as to increase TF-cooperativity potency (i.e., ). Specifically, we calculate the nominal promoter unbinding rates via the formula . Where, for a given gene-TF pair: ; kb is its nominal promoter binding rate; q is its maximal occupancy.
@@
The exploration of the immense parameter space of a mechanistic model is a demanding computational task, especially for biophysical problems dealing with spatial-stochastic system representations [147]. Several classical inference/optimization methods such as heuristic tuning, stochastic gradient descent, simulated annealing, and approximate Bayesian computation (ABC) have been traditionally used for finding reasonable parameter sets of biophysical mechanistic models [116, 27, 50, 133]. Nonetheless, all of these aforementioned methods are often computationally inefficient because they require many expensive simulations for properly scanning the model parameter space, and they commonly lack powerful parameter-space interpolation strategies. As a consequence, these inference methods become prohibitively ineffective when they are applied to high-dimensional spatial-stochastic models, which are nowadays usually employed to represent complex biological systems.
@@
Biophysically-realistic mechanistic modeling is integral to achieving a quantitative understanding of the behavior of complex biological systems. These biophysical representations are fundamentally generative models: they aim at providing a mechanistic description of the underlying biological phenomenon, enabling the generation of temporally-faithful synthetic trajectories of the given modeled dynamics, which potentially resemble empirical findings or target observations. As such, these models require a rich collection of experimental data detailing the principal mechanisms and physical processes triggering the phenomenon under study.
@@
However, the SBI framework has recently emerged as an influential technique to tackle this inferential challenge [27]. SBI does not attempt to retrieve the single best parameter set. Instead, SBI helps to discover high-probability parameter space regions capable of explaining the target observations, which explicitly quantifies parameter uncertainty [136].
@@
In our case, to overcome this challenging inferential task and inspired by the simulation-based inference (SBI) framework, we perform a novel AI-powered parameter estimation approach which blends itself with various classical ML techniques. At its core, this novel inference approach directly exploits simulation data for exploring parameter space, yet it indirectly relies on primarily qualitative data to inform/constrain the parameter values; the simulation data is accordingly generated via the proposed spatial-stochastic model of mouse ICM lineage differentiation. Our model parameter estimation scheme is thus guided by the SBI paradigm [27], and its centerpiece is the sequential neural posterior estimation (SNPE) algorithm [59, 32].
@@
A central idea of biochemical network model building is to reproduce some particularly interesting behaviors, phenotypes, or empirical observations [133]. Nevertheless, because of highly-nonlinear interactions among genetic regulatory components, a single set of experimental data by itself can not uncover the full functional dynamics of such biological systems [116]. It is hence normally necessary to consolidate several datasets coming from multiple related systems under diverse experimental conditions; for this reason, coherent and viable data analysis becomes a considerably difficult endeavor on its own. What is more, these quantitative observations often do not provide enough details for spatial-stochastic modeling, and they are typically incomplete: from the practical point of view, it is extremely challenging to provide simultaneous measurements for all the key elements dictating the underlying multiscale dynamics.
@@ Acknowledgments
The successful completion of this research project owes much to the collaboration and support of esteemed colleagues and collaborators. We extend our deepest gratitude to Sabine Fischer, Tim Liebisch, Franziska Matthäus, and Simon Schardt for their pivotal roles in fostering insightful discussions and providing constructive feedback that significantly elevated the quality of our research.
@@
Furthermore, we would like to acknowledge the CMMS project and FIAS for their financial support, which has been instrumental in facilitating the progress of this research. Their commitment to advancing scientific endeavors has been a cornerstone in our pursuit of knowledge.
Lastly, we extend our thanks to all individuals who have contributed, directly or indirectly, to the success of this project. Your collective efforts have undeniably contributed to the depth of our scientific inquiry.

### 13 Mar 2024 v2

@@ Discussion
The first cell-fate decision happens between E2.5 and E3.0 (from 8- to 32-cell stage): cells acquire TE or ICM identities. The second cell-fate decision happens at the ICM between E3.0 and E4.0 (from 32- to 128-cell stage). From E4.0 to 4.5, EPI and PRE populations spatially separate.
While it is customary to define the blastocyst-formation period between E3.0 and E4.5 [11, 25], these boundaries are ultimately arbitrary as development occurs in a continuum and diverse experimental arrangements/conditions are in use between distinct labs. Moreover, ICM cells adopt their next identities asynchronously as the blastocyst forms [84, 12, 24]. Together with these aspects, it is also commonly accepted that cells are already coexpressing Nanog- and Gata6-related factors at around E2.75 [135, 19], plus Fgf4 expression is already perceptible at around E3.25 [11, 25]. For these reasons, our standard model simulations target a time-window of 48 hours (E2.75-E4.75); this range allows us to circumvent potential discrepancies among timing annotations and keep a temporally faithful description of the biological system under study.
@@
The core GRN motif is exclusively comprised by Nanog- and Gata6-related elements. To be more precise, all their directly related species and interactions. The rest of the full GRN is built around the core motif, thus consolidating the remaining elements and their collective effects on the dynamics of the two main players.
Importantly, we have arranged all cell-scale reaction events into several groups as follows: summary of gene expression dynamics; promoter binding and unbinding; mRNA synthesis and degradation; protein synthesis and degradation; FGFR activation and inactivation; ERK activation and inactivation; NANOG phosphorylation and dephosphorylation. In that regard, we report all the particular interactions implemented by our simulator and their respective literature sources.
@@
We model TF cooperativity by expressly tuning the promoter unbinding rates. This rate tuning influences the promoter regulation model to mimic a Hill-function-like (nonlinear) transcriptional response. The usage of the Hill function is a staple of phenomenological modelling, however it is incompatible with mechanistic modelling; directly using a Hill equation as a reaction propensity function ignores the non-instantaneous (stochastic) nature of delays between biochemical events and introduces several other simulation artifacts [150, 151, 152].
We use the concepts of a half-saturation constant and a cooperativity coefficient to perform promoter unbinding rate tuning; accordingly, both quantities are incorporated into elementary reactions to describe promoter unbinding dynamics. This half-saturation constant (hact or hrep depending on TF role) is a free model parameter which dictates the threshold of TF copies needed for reaching 50% of negative or positive gene transcriptional control; consequently, each gene-TF pair requires its own separate half-saturation threshold. This cooperativity coefficient (kcoop) is an auxiliary variable which adjusts the strength of mutual influence among TF copies; we arbitrarily defined it as to increase TF-cooperativity potency (i.e., ). Specifically, we calculate the nominal promoter unbinding rates via the formula . Where, for a given gene-TF pair: ; kb is its nominal promoter binding rate; q is its maximal occupancy.
@@
The exploration of the immense parameter space of a mechanistic model is a demanding computational task, especially for biophysical problems dealing with spatial-stochastic system representations [108]. Several classical inference/optimization methods such as heuristic tuning, stochastic gradient descent, simulated annealing, and approximate Bayesian computation (ABC) have been traditionally used for finding reasonable parameter sets of biophysical mechanistic models [81, 74, 82, 80, 176]. Nonetheless, all of these aforementioned methods are often computationally inefficient because they require many expensive simulations for properly scanning the model parameter space, and they commonly lack powerful parameter-space interpolation strategies. As a consequence, these inference methods become prohibitively ineffective when they are applied to high-dimensional spatial-stochastic models, which are nowadays usually employed to represent complex biological systems.
@@
Biophysically-realistic mechanistic modeling is integral to achieving a quantitative understanding of the behavior of complex biological systems. These biophysical representations are fundamentally generative models: they aim at providing a mechanistic description of the underlying biological phenomenon, enabling the generation of temporally-faithful synthetic trajectories of the given modeled dynamics, which potentially resemble empirical findings or target observations. As such, these models require a rich collection of experimental data detailing the principal mechanisms and physical processes triggering the phenomenon under study [110].
@@
However, an AI-powered SBI framework has recently emerged as an influential technique to tackle any likelihood-free inferential challenge [74]. In general, SBI does not attempt to retrieve the single best parameter set. Instead, SBI helps to discover high-probability parameter space regions capable of explaining the target observations, while potentially facilitating the quantification of parameter uncertainty [109].
@@
In this work, to overcome our challenging inferential task and inspired by the simulation-based inference (SBI) framework, we exploit this novel AI-powered parameter estimation approach and combine it with various classical ML techniques. At its core, our inference approach directly exploits simulation data for exploring parameter space, yet it indirectly relies on primarily qualitative data to inform/constrain the parameter values; the simulation data is accordingly generated via the proposed spatial-stochastic model of mouse ICM lineage differentiation. Our model parameter estimation scheme is thus guided by the SBI paradigm [74], and its centerpiece is the sequential neural posterior estimation (SNPE) algorithm [71, 72].
@@
A central idea of biochemical network model building is to reproduce some particularly interesting behaviors, phenotypes, or empirical observations [80]. Nevertheless, because of highly-nonlinear interactions among genetic regulatory components, a single set of experimental data by itself can not uncover the full functional dynamics of such biological systems [81]. It is hence normally necessary to consolidate several datasets coming from multiple related systems under diverse experimental conditions; for this reason, coherent and viable data analysis becomes a considerably difficult endeavor on its own.
What is more, these quantitative observations often do not provide enough details for spatial-stochastic modeling, and they are typically incomplete: from the practical point of view, it is extremely challenging to provide simultaneous measurements for all the key elements dictating the underlying multiscale dynamics.
@@ Acknowledgments
The successful completion of this research project owes much to the collaboration and support of esteemed colleagues and collaborators. We extend our deepest gratitude to Sabine Fischer, Tim Liebisch, Franziska Matthäus, and Simon Schardt for their pivotal roles in fostering insightful discussions and providing constructive feedback.
@@
Furthermore, we would like to acknowledge the CMMS project and FIAS for their financial support, which has been instrumental in facilitating the progress of this research. We also thank the Center for Scientific Computing (CSC) at Goethe University Frankfurt for granting us access to the Goethe-HLR cluster.